OpenStudy (anonymous):

Identify the correct slope and y-intercept of the equation 2x + 3y = 6 slope = -2; y-intercept = (0, 6) slope = -three halves; y-intercept = (0, 3) slope = -2; y-intercept = (0, 3) slope = -two thirds; y-intercept = (0, 2)

OpenStudy (anonymous):

last one

OpenStudy (katrinakaif):

Well turn this into a standard y= mx + b format: so 2x + 3y = 6 3y = -2x + 6 y = -2/3x + 6/3 y = -2/3x + 2 ---> y = mx + b M is the slope. and here it is 2/3.